OpenStudy (anonymous):

At time t=0, a particle is located at the point (1,2,3). It travels in a straight line to the point (4,1,4), has speed 2 at (1,2,3) and constant acceleration 3i-j+k. FInd position vector r(t) at t. I got v(t) = 3ti -tj +tk +C, and |v(0)| =2, the direction would be 3i-j+k. I'm stuck at this point. Help!

OpenStudy (abb0t):

Given r(0) = (1, 2, 3) and speed |v(0)| = 2 at \[a(t) = \frac{ dv }{ dt } = 3î - j + k\] integrate on both sides. where the unit vector is \[\frac{ u }{ |u| }\] and \[|u| = \sqrt{a^2 + b^2 +c^2}\] your answer should be \[\sqrt{11}\] for this part.

OpenStudy (anonymous):

I'm confused. Is sqrt(11) the length of acceleration vector ? I think the question ask for r(t)

OpenStudy (abb0t):

This you'r unit vector(s) are: \[(\frac{ 3 }{ \sqrt{11} }), (-\frac{ 1 }{ \sqrt{11} }), (\frac{ 1 }{ \sqrt{11} }) \] now unit vector multiplied by speed gives \[\frac{ u }{ |u| }\times v\] which is the same thing as v(0).

OpenStudy (abb0t):

r(t) is your position vector. And it will be a function in terms of t.

OpenStudy (abb0t):

You have to get your velocity function first so that you can integrate to get your position function, which is why I am doing this.

OpenStudy (abb0t):

Can you take it from here or did you still need more help?

OpenStudy (anonymous):

So, v(t) = 3ti -tj+tk +C

OpenStudy (abb0t):

Correct.

OpenStudy (anonymous):

i think I'm stuck right here :D

OpenStudy (abb0t):

Well, v(0) = A (I used a different constant. A for acceleration) therefore, your velocity vector is: \[v(t) = (3t)i - (t)j + (t)k + v(0) \] and you already have v(0) which I gave to you before. it was your \[\frac{ u }{ |u| }\times v\]

OpenStudy (abb0t):

Just a hint, your x component (your "i" vector) should start with \[v(t) = (3t + \frac{ 6 }{ \sqrt{11} })i - ...\]

OpenStudy (abb0t):

Once you have your v(t) simply \[\int\limits v(t) dt\]

OpenStudy (anonymous):

I know this part, but still thinking about how u get r(t) unit vectors

OpenStudy (abb0t):

\[\int\limits v(t) dt = r(t) + B\]

OpenStudy (anonymous):

oh I got it. you indicate u for acceleration, that's why i'm confused. Thanks!.

OpenStudy (anonymous):

v(t) = 3/sqrt(11)ti - 1/sqrt(11)tj +1/sqrt(11)tk

OpenStudy (abb0t):

Your answer should be: \[r(t) = (\frac{ 3t^2 }{ 2 }+\frac{ 6 }{ \sqrt{11}}t+1)i - (\frac{ t^2 }{ 2 }+\frac{ 2 }{ \sqrt{11}}t-2)j+(\frac{ t^2 }{ 2 }+\frac{ 2 }{ \sqrt{11}}t+3)k\]